OpenStudy (anonymous):

Simplify: (6b^-1*6a^4b^6)/(-3a^0*5a^4b^0)

OpenStudy (jmark):

(6b^-1*6a^4b^6)/(-3a^0*5a^4b^0) = -(6*6a^4b^6)/( b*3(1)*5a^4(1)) = -(36*a^4*b^6)/( 15*a^4*b) divide the common term = - (12*b^5)/( 5) ICSE class 8 syllabus - http://icse.edurite.com/icse-syllabus/icse-syllabus-for-class-8.html